# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ Appendix H Helper Lemmas
Proof.
Using a result from Li [2010], we write
@@
Meanwhile, a simple but tight bound Alzer [2001] for the denominator is
@@
Lemma H.2 (Adapted from Appendix G.4 of Agarwal et al. [2022]).
Let $\mathbf{Y}=\mathbf{M}+\mathbf{E}\in\mathbb{R}^{m\times n}$ with $\mathsf{rank}(\mathbf{M})=d$. Let $\mathbf{U}_{d}\mathbf{\Sigma}_{d}\mathbf{V}_{d}^{\top}$ and $\hat{\mathbf{U}}_{d}\hat{\mathbf{\Sigma}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ denote the top $d$ singular components of the SVD of $\mathbf{M}$ and $\mathbf{Y}$ respectively. Then, the truncated SVD estimator $\hat{\mathbf{M}}=\hat{\mathbf{U}}_{d}\hat{\mathbf{\Sigma}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ is such that for all $j\in[m]$,
$\left\lVert\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}\right\rVert_{2}^{2}\leq 2\cdot\left\lparen\frac{\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}^{2}}{\sigma_{d}^{2}(\mathbf{M})}\left\lparen\left\lVert\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{M}_{j}\right\rVert_{2}^{2}\right\rparen+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen.$ ) . (157)
Proof.
@@
$\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}$ $=(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{Y}_{j}-\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j})+(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j}-\mathbf{M}_{j})$ (158)
$=\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}+(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{I}_{n})~{}\mathbf{M}_{j}.$ (159)
@@
$\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}$ $\leq 2\cdot\left\lparen\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen$ ) (161)
@@
Since $\mathsf{rank}(\mathbf{M})=d$, the Wedin $\sin{\Theta}$ Theorem Wedin [1972] guarantees that
@@
Combining (160), (164), and (167) we obtain the desired bound. ∎
Lemma H.3 (Proposition 2.4 of Rudelson and Vershynin [2010]).
@@
Lemma H.4 (Matrix Bernstein inequality).
@@
$\leq\begin{cases}d\exp{\left(-t^{2}/4\sigma^{2}\right)}&\quad\text{for $t\leq\sigma^{2}/R$}\\ d\exp{\left(-t^{2}/4R\right)}&\quad\text{for $t\geq\sigma^{2}/R$}\end{cases}.$
Lemma H.5 (Theorem 5.39 of Vershynin [2010]).
@@
$\sqrt{n}-C\sqrt{d}-t\leq\sigma_{\min}(\mathbf{X})\leq\sigma_{\max}(\mathbf{X})\leq\sqrt{n}+C\sqrt{d}+t,$
@@
Lemma H.6 (Mean of independent sub-gaussian random variables).
@@
Lemma H.7 (Norm of sub-gaussian random vector).
@@
Proof.
Since $\mathbf{x}$ is $\sigma^{2}$-sub-gaussian, it is $(n\sigma^{2})$-norm-sub-gaussian as defined Jin et al. [2019]. Using Definition 3 and Lemma 1 of Jin et al. [2019], the result immediately follows. ∎
Lemma H.8 (Lemma 12 of Shah et al. [2020]).
Let $M=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. If $\mathsf{rank}(A)=\mathsf{rank}(M)$, then $D=CA^{\dagger}B$.
Lemma H.9 (Theorem 3.3 of Stewart [1977]).

### v3

@@ Appendix D Helper Lemmas
Proof of Lemma 10.
Using a result from [27], we write
@@
Meanwhile, a simple but tight bound [7] for the denominator is
@@
Lemma 11 (Adapted from Appendix G.4 of [2]).
Let $\mathbf{Y}=\mathbf{M}+\mathbf{E}\in\mathbb{R}^{m\times n}$ with $\mathrm{rank}(\mathbf{M})=d$. Let $\mathbf{U}_{d}\mathbf{\Sigma}_{d}\mathbf{V}_{d}^{\top}$ and $\hat{\mathbf{U}}_{d}\hat{\mathbf{\Sigma}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ denote the top $d$ singular components of the SVD of $\mathbf{M}$ and $\mathbf{Y}$ respectively. Then, the truncated SVD estimator $\hat{\mathbf{M}}=\hat{\mathbf{U}}_{d}\hat{\mathbf{\Sigma}}_{d}\hat{\mathbf{V}}_{d}^{\top}$ is such that for all $j\in[m]$,
$\left\lVert\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}\right\rVert_{2}^{2}\leq$
$\qquad 2\cdot\left\lparen\frac{\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}^{2}}{\sigma_{d}^{2}(\mathbf{M})}\left\lparen\left\lVert\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{M}_{j}\right\rVert_{2}^{2}\right\rparen+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen.$ ) .
Proof of Lemma 11.
@@
$\hat{\mathbf{M}}_{j}-\mathbf{M}_{j}$ (103)
$\quad=(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{Y}_{j}-\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j})+(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{M}_{j}-\mathbf{M}_{j})$ (104)
$\quad=\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}+(\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}-\mathbf{I}_{n})~{}\mathbf{M}_{j}.$ (105)
@@
$\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}$ (107)
$\ \ \leq 2\cdot\left\lparen\left\lVert\hat{\mathbf{V}}_{d}\hat{\mathbf{V}}_{d}^{\top}\mathbf{E}_{j}-\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}+\left\lVert\mathbf{V}_{d}\mathbf{V}_{d}^{\top}\mathbf{E}_{j}\right\rVert_{2}^{2}\right\rparen$ ) (108)
@@
Since $\mathrm{rank}(\mathbf{M})=d$, the Wedin $\sin{\Theta}$ Theorem [42] guarantees that
@@
Combining (106), (111), and (114) we obtain the desired bound.
Lemma 12 (Proposition 2.4 of [33]).
@@
Lemma 13 (Matrix Bernstein inequality).
@@
Lemma 14 (Theorem 5.39 of [41]).
@@
$\forall i\in[d],\ \sqrt{n}-C\sqrt{d}-t\leq\sigma_{i}(\mathbf{X})\leq\sqrt{n}+C\sqrt{d}+t,$
@@
Lemma 15 (Mean of independent sub-gaussian random variables).
@@
Lemma 16 (Norm of sub-gaussian random vector).
@@
The proof of Lemma 16 is immediate. Since $\mathbf{x}$ is $\sigma^{2}$-sub-gaussian, it is $(n\sigma^{2})$-norm-sub-gaussian as defined [23]. Using Definition 3 and Lemma 1 of [23], the result immediately follows.
Lemma 17 (Lemma 12 of [37]).
Let $M=\begin{bmatrix}A&B\\ C&D\end{bmatrix}$. If $\mathrm{rank}(A)=\mathrm{rank}(M)$, then $D=CA^{\dagger}B$.
Lemma 18 (Theorem 3.3 of [38]).
@@
Lemma 19 (Adapted from Proposition 13 of [37] ).
Let $\mathbf{H}$ be an $|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|$ dimensional matrix with entries are in $[-dB^{2},dB^{2}]$, and satisfying $\mathrm{rank}(\mathbf{H})=d$. Let $\overline{\mathbf{H}}$ be an alternative $|\mathcal{U}^{\star}|\times|\mathcal{V}^{\star}|$ dimensional matrix. Now, for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$, assume that there exists $\mathcal{I}\subseteq\mathcal{U}^{\star}$ and $\mathcal{J}\subseteq\mathcal{V}^{\star}$, such that $|\mathcal{I}|=|\mathcal{J}|=d$, and for all $k\in\mathcal{I},\ell\in\mathcal{J}$,
$|\mathbf{H}_{i\ell}-\overline{\mathbf{H}}_{i\ell}|\leq\epsilon,\ |\mathbf{H}_{k\ell}-\overline{\mathbf{H}}_{k\ell}|\leq\epsilon,\ |\mathbf{H}_{kj}-\overline{\mathbf{H}}_{kj}|\leq\epsilon,$ (121)
it holds that
Proof of Lemma 19.
First, observe that for any $i\in\mathcal{U}^{\star},j\in\mathcal{V}^{\star}$, we have by Lemma 17 that
as $\mathbf{H}_{\mathcal{I},\mathcal{J}}$ and $\mathbf{H}$ both have rank $d$.
$\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}$
$\qquad=\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j}-\mathbf{H}_{i,\mathcal{J}}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}$ (125)
$\qquad=\overline{\mathbf{H}}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\overline{\mathbf{H}}_{\mathcal{I},j}-\mathbf{H}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}$
$\qquad\ \ \ +\mathbf{H}_{i,\mathcal{J}}\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}-\mathbf{H}_{i,\mathcal{J}}\left(\mathbf{H}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\mathbf{H}_{\mathcal{I},j}$ (126)
$\qquad=\mathrm{Tr}\left(\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\left(\overline{\mathbf{H}}_{\mathcal{I},j}\overline{\mathbf{H}}_{i,\mathcal{J}}-\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right)\right)$
$\qquad\ \ \ +\mathrm{Tr}\left(\left((\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right)\left(\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right)\right)$ (127)
Since $\lvert\mathrm{Tr}(AB)\rvert\leq\sqrt{\mathrm{rank}(B)}\left\lVert A\right\rVert_{\operatorname{op}}\left\lVert B\right\rVert_{F}$, we obtain
$\left\lvert\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}\right\rvert$ (128)
$\quad\leq\sqrt{2}\left\lVert\left(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}}\right)^{\dagger}\right\rVert_{\operatorname{op}}\left\lVert\overline{\mathbf{H}}_{\mathcal{I},j}\overline{\mathbf{H}}_{i,\mathcal{J}}-{\mathbf{H}}_{\mathcal{I},j}{\mathbf{H}}_{i,\mathcal{J}}\right\rVert_{F}$ (129)
$\quad+\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}\left\lVert\mathbf{H}_{\mathcal{I},j}\mathbf{H}_{i,\mathcal{J}}\right\rVert_{F}.$ (130)
$\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}$ $=\frac{1}{\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})}$
$\leq\frac{1}{\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})-2\epsilon d}$
$\leq\frac{2}{\sigma_{d}(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})}.$
To bound the second part of the expression, note $\left\lVert\overline{\mathbf{H}}_{\mathcal{I},j}\overline{\mathbf{H}}_{i,\mathcal{J}}-{\mathbf{H}}_{\mathcal{I},j}{\mathbf{H}}_{i,\mathcal{J}}\right\rVert_{F}\leq 2d(2dB^{2}\epsilon+\epsilon^{2})$ because $\left\lVert\overline{\mathbf{H}}_{i^{\prime},j}\overline{\mathbf{H}}_{i,j^{\prime}}-{\mathbf{H}}_{i^{\prime},j}{\mathbf{H}}_{i,j^{\prime}}\right\rVert_{F}\leq 2dB^{2}\epsilon+\epsilon^{2}$ for all $(i^{\prime},j^{\prime})\in\mathcal{U}^{\star}\times\mathcal{V}^{\star}$. Now, to upper bound (130), we first upper bound $\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}$ by using Lemma 18 to note
$\left\lVert(\overline{\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}-({\mathbf{H}}_{\mathcal{I},\mathcal{J}})^{\dagger}\right\rVert_{\operatorname{op}}\leq\frac{1+\sqrt{5}}{2}\cdot\frac{8\epsilon d.}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}$ (131)
$\left\lvert\widehat{\mathbf{H}}_{i,j}-\mathbf{H}_{i,j}\right\rvert$ (132)
$\qquad\leq\left(\frac{4\sqrt{2}(d^{2}B^{2}+d\epsilon)}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\epsilon$ (133)
$\qquad\leq\left(\sqrt{2}+\frac{4\sqrt{2}d^{2}B^{2}}{\sigma_{d}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}+\frac{2(1+\sqrt{5})4d^{3}B^{2}}{\sigma_{d}^{2}(\mathbf{H}_{\mathcal{I},\mathcal{J}})}\right)\epsilon$ (134)